# Robust Implicit Regularization via Weight Normalization

### v3

@@ 3.2 Invariants of the flow
Since $\gamma$ defined in (42) is both bounded above and below, it suffices to prove the case for ${\bf x}$. We will prove the statement by contradiction. Suppose $x_{j}$ is not bounded away from zero. Then, there exists a sequence $\{t_{k}\}_{k\in\mathbb{N}}$ such that $\lim_{k\to\infty}x_{j}(t_{k})=0$. By assumption there exists ${\bf v}>0$ such that ${\bf A}{\bf v}=0$. Then for any ${\bf z}$ such that $(I-\mathcal{P}_{{\bf A}}){\bf z}=0$, it holds
@@
Lemma 3.11 (Uniqueness).
$\lim_{t\to\infty}{\bf A}[\tilde{{\bf x}}^{(1)}(t)-\tilde{{\bf x}}^{(2)}(t)]=0$ (43)
and
$\begin{cases}\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)[\log(\tilde{{\bf x}}^{(1)}(t))-\log(\tilde{{\bf x}}^{(2)}(t))]=0&\text{if }L=2,\\ \lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)[(\tilde{{\bf x}}^{(1)}(t))^{\odot\frac{2}{L}-1}-(\tilde{{\bf x}}^{(2)}(t))^{\odot\frac{2}{L}-1}]=0&\text{if }L\neq 2,\end{cases}$ (44)
then $\lim_{t\to\infty}\tilde{{\bf x}}^{(1)}(t)-\tilde{{\bf x}}^{(2)}(t)=0$.
Proof.
@@
$\geq\bm{\Delta}\odot\log({\bf 1}+\bm{\Delta}\odot(\tilde{{\bf x}}^{(1)})^{\odot-1}).$ (45)
Since $\bm{\xi}\geq 0$ and $\langle\bm{\xi}(t),{\bf 1}\rangle$ converges to zero as $t\to\infty$, we can deduce that $\bm{\xi}$ converges to zero as $t\to\infty$. Together with (45) and the assumption that $\tilde{{\bf x}}^{(1)}$ is uniformly bounded above, we conclude that $\bm{\Delta}$ must also converges to zero. Since $\bm{\Delta}$ converges to zero, our conclusion follows.
@@
Proof of Theorem 2.2.
By Lemma 3.9, we also have
Since we can choose $r_{0}$ but not $r_{\infty}$, we will examine the effect of $r_{\infty}$ while fixing $r_{0}$. For now, let us assume that $r_{\infty}\gg r_{0}$. The partial derivative of $\gamma$ is given by
This implies that as long as ${\tilde{\eta}}\leq r_{0}^{2}$, we have
$\gamma(r_{0},r_{\infty})=\gamma(r_{0},r_{0})+\int_{r_{0}}^{r_{\infty}}\partial_{r}\gamma(r_{0},r)dr\leq\gamma(r_{0},r_{0})=1.$
In fact, for any $r\in[r_{0},r_{\infty}]$, $\gamma(r_{0},r_{\infty})\leq\gamma(r_{0},r)$. Hence to upper bound $\gamma(r_{0},r_{\infty})$, it suffices to lower bound $r_{\infty}$. By assumption ${\bf A}\tilde{{\bf x}}_{\infty}={\bf b}$ and hence $\|\tilde{{\bf x}}_{\infty}\|_{2}\geq\|{\bf A}^{\dagger}{\bf b}\|_{2}$. Consequently
$r_{\infty}$ $=\|\tilde{{\bf x}}_{\infty}^{\odot 1/L}\|_{2}=\|\tilde{{\bf x}}_{\infty}\|_{2/L}^{1/L}$
$\geq\|\tilde{{\bf x}}_{\infty}\|_{2}^{1/L}\geq\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}.$
This also proves that $r_{\infty}\geq r_{0}$ and the proof is complete. ∎
3.3 Convergence for L=1,2
Lemma 3.12 (Boundedness).
Let $L=1,2$ and ${\tilde{\eta}}>0$. Suppose $(r,{\bf u})$ follow the dynamics in (13) and (14) with $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$, $r_{0},{\bf u}_{0}>0$. If all entries of ${\bf u}$ are bounded away from zero, then $r$ is uniformly upper bounded.
Proof.
@@ 3.3 Convergence for L=1,2
$\|\mathcal{P}_{{\bf A}}(r{\bf u})^{\odot L}\|_{\infty}\leq B_{1}$ (46)
@@
$(I-\mathcal{P}_{{\bf A}})\cdot\begin{cases}\log(\widetilde{r}{\bf u})&\text{if}L=2\\ (\widetilde{r}{\bf u})^{\odot 2-L}&\text{if }L\neq 2\end{cases}$ (47)
@@
$\langle(r{\bf u})^{\odot 2},\log(\widetilde{r}{\bf u})\rangle\geq r^{2}\left(\log(\widetilde{r})-\frac{N}{2e}\right).$ (48)
@@
Note that $B_{1}\left\|\mathcal{P}_{{\bf A}}\log\left({\bf u}\right)\right\|_{1}$ is upper bounded by some constant $C$ because the entries of ${\bf u}$ are bounded both above and below. Combining this with (48), we have
@@
Theorem 3.13 (Theorem 4 in [5]).
If $\mathcal{L}:\mathbb{R}^{N}\to\mathbb{R}$ is analytic and the curve $t\mapsto{\bf x}(t)\in\mathbb{R}^{N}$, $t\in[0,\infty)$ is bounded and a solution of the gradient flow equation $\partial_{t}{\bf x}=\nabla\mathcal{L}({\bf x})$, then ${\bf x}$ converges to a critical point of $\mathcal{L}$ as $t\to\infty$.
Proof of Theorem 2.3.
3.4 An example of time-dependent learning rate
In this section we study a particular example of time-dependent learning rate, given by $(\eta_{r},\eta_{\bf u})=(r^{2},1)$. Note that instead of a constant in time, $\eta_{r}=r^{2}(t)$ is a function depends on time. In this case the dynamics is greatly simplified and is similar to gradient flow without normalization (11). Such simplification allows us to analyze the dynamics based on established methods, such as the argument with Bregman divergence in [8], and completely bypass the need of invariants and uniqueness results proved in Section 3.2.
However, gradient flow under this particular choice of learning rate ($(\eta_{r},\eta_{\bf u})=(r^{2},1)$) no longer exhibits the magnification effect as in the constant rate case ($(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$), and hence does not yield better bounds than the ones in previous works (8). It is nevertheless remarkable that the dynamics with certain choices of learning rate can be so different from the one with time-independent learning rate.
Lemma 3.14 (Dynamics reduction).
Suppose $r$ and ${\bf u}$ follow the gradient flow in (13) and (14) with $\|{\bf u}(0)\|=1$. Let ${\bf x}:=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Then
$\partial_{t}{\bf x}=-\left(\eta_{r}\frac{{\bf x}{\bf x}^{\top}}{\|{\bf x}\|_{2}^{2}}+\eta_{\bf u}(\|{\bf x}\|_{2}^{2}I-{\bf x}{\bf x}^{\top})\right)\nabla\mathcal{L}\left({\bf x}\right).$ (49)
Proof.
By (23) and (24),
@@ 3.4 An example of time-dependent learning rate
and hence (49) holds. ∎
Observe that if $(\eta_{r},\eta_{\bf u})=(r^{2},1)$, then (49) is greatly simplified as stated next.
Corollary 3.15 (Dynamic reduction, time-dependent learning rate).
$\partial_{t}{\bf x}=-\|{\bf x}\|_{2}^{2}\cdot\nabla\mathcal{L}\left({\bf x}\right).$ (50)
Proof.
The statement follows from Lemma 3.14 by substituting $(\eta_{r},\eta_{\bf u})=(r^{2},1)$ into (49). ∎
Note that without the additional factor $\|{\bf x}\|_{2}^{2}$, we are back to the setting that does not include weight normalization at all (3). Although the extra factor changes in time, since it is only a scalar as opposed to a vector or a matrix, it is possible to apply the same proof strategy as for gradient flow without normalization as in [8].
Definition 3.16 (Bregman Divergence).
$D_{F}(p,q)=F(p)-F(q)-\langle\nabla F(q),p-q\rangle.$ (51)
Lemma 3.17 ([6]).
Lemma 3.18 (Non-increasing Bregman Divergence).
$\partial_{t}D_{F}({\bf z},\tilde{{\bf x}}(t))=-2L\|{\bf x}(t)\|_{2}^{2}\cdot\mathcal{L}({\bf x}(t))$ (52)
$F(\tilde{{\bf x}})=\begin{cases}\frac{1}{2}\langle\tilde{{\bf x}}\odot\log(\tilde{{\bf x}})-\tilde{{\bf x}},\bm{1}\rangle&\text{if }L=2\\ \frac{L}{2(2-L)}\langle\tilde{{\bf x}}^{\odot\frac{2}{L}},\bm{1}\rangle&\text{if }L>2.\end{cases}$ (53)
Lemma 3.19 (Convergence, time-dependent learning rate).
Let $L\in\mathbb{N}$, $L\geq 2$, $(\eta_{r},\eta_{\bf u})=(r^{2},1)$, $\|{\bf u}(0)\|=1$ and $r(0)>0$. Denote ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Suppose that ${\bf b}$ is not identically zero and the solution set $S_{+}=\{{\bf z}\geq 0:{\bf A}{\bf z}={\bf b}\}$ is non-empty. Then
$\lim_{t\to\infty}\mathcal{L}({\bf x}(t))=0.$
Proof.
@@
Let ${\bf z}\in S_{+}$. Since $r(0)>0$ and ${\bf u}(0)\geq 0$, we have ${\bf x}\geq 0$ for all $t\geq 0$ by Lemma 3.3. Hence $D_{F}({\bf z},\tilde{{\bf x}}(t))$ is well-defined. By Lemma 3.18, $D_{F}({\bf z},\tilde{{\bf x}}(t))$ is non-increasing in $t$ and hence bounded above by $D_{F}({\bf z},\tilde{{\bf x}}_{0})$. By Lemma 3.17, $D_{F}$ is non-negative and hence bounded below. Therefore by (52),
@@
Since $D_{F}({\bf z},\tilde{{\bf x}}_{0})<\infty$, by (54) $\|{\bf x}(t)\|_{2}$ cannot be bounded away from zero. This implies that there exists an increasing sequence $\{t_{k}\}_{k\in[\mathbb{N}]}$ such that $\lim_{k\to\infty}\|{\bf x}(t_{k})\|_{2}=0$. Together with the fact that ${\bf x}$ is non-negative, we get that for all $n\in[N]$, $\lim_{k\to\infty}x_{n}(t_{k})=0$, and consequently $\lim_{k\to\infty}\tilde{x}_{n}(t_{k})=0$.

### v4

@@ 3.2 Invariants of the flow
See Appendix A. ∎
Lemma 3.11 (Uniqueness).
$\lim_{t\to\infty}{\bf A}[\tilde{{\bf x}}^{(1)}(t)-\tilde{{\bf x}}^{(2)}(t)]=0$ (3.17)
and
$\begin{cases}\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)[\log(\tilde{{\bf x}}^{(1)}(t))-\log(\tilde{{\bf x}}^{(2)}(t))]=0&\text{if }L=2,\\ \lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)[(\tilde{{\bf x}}^{(1)}(t))^{\odot\frac{2}{L}-1}-(\tilde{{\bf x}}^{(2)}(t))^{\odot\frac{2}{L}-1}]=0&\text{if }L\neq 2,\end{cases}$ (3.18)
then $\lim_{t\to\infty}\tilde{{\bf x}}^{(1)}(t)-\tilde{{\bf x}}^{(2)}(t)=0$.
Proof.
See Appendix A. ∎
Proof of Theorem 2.2.
By Lemma 3.9, we also have
We now examine the relation between $r_{0}$, $r_{\infty}$, and $\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}$. Because ${\bf A}^{\dagger}{\bf b}\in\operatorname*{arg\,min}_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{2}$ and ${\bf A}\tilde{{\bf x}}_{\infty}={\bf b}$, we have
$r_{\infty}=\|\tilde{{\bf x}}_{\infty}^{\odot 1/L}\|_{2}=\|\tilde{{\bf x}}_{\infty}\|_{2/L}^{1/L}\geq\|\tilde{{\bf x}}_{\infty}\|_{2}^{1/L}\geq\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}\geq r_{0}.$ (3.19)
Recall that
$\gamma(r_{0},r)=\frac{r}{r_{0}}\exp\left(\frac{r_{0}^{2}-r^{2}}{2{\tilde{\eta}}}\right).$
The partial derivative of $\gamma$ is given by
Note that $\partial_{r}\gamma(r_{0},r)\leq 0$ for $r\geq\sqrt{{\tilde{\eta}}}$. By (3.19), we have $r_{\infty}\geq\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}\geq r_{0}\geq\sqrt{{\tilde{\eta}}}$ and hence
$\gamma(r_{0},r_{\infty})\leq\gamma(r_{0},\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L})\leq\gamma(r_{0},r_{0})=1.$
Thus $\gamma(r_{0},\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L})$ is an upper bound of $\gamma(r_{0},r_{\infty})$. Because $\rho=\gamma^{-1}$ represents the improvement under weight normalization (larger $\rho$ is better), $\gamma(r_{0},\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L})^{-1}$ is a lower bound, or a minimal guarantee, for the improvement we will get. This completes the proof. ∎
3.3 Convergence for L=1,2
Lemma 3.12 (Boundedness).
Let $L=1,2$ and ${\tilde{\eta}}>0$. Suppose $(r,{\bf u})$ follow the dynamics in (1.13) and (1.14) with $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$, $r_{0},{\bf u}_{0}>0$. If all entries of ${\bf u}$ are bounded away from zero, then $r$ is uniformly upper bounded.
Proof.
See Appendix A. ∎
Theorem 3.13 (Theorem 4 in [5]).
If $\mathcal{L}:\mathbb{R}^{N}\to\mathbb{R}$ is analytic and the curve $t\mapsto{\bf x}(t)\in\mathbb{R}^{N}$, $t\in[0,\infty)$ is bounded and a solution of the gradient flow equation $\partial_{t}{\bf x}=-\nabla\mathcal{L}({\bf x})$, then ${\bf x}$ converges to a critical point of $\mathcal{L}$ as $t\to\infty$.
Proof of Theorem 2.3.
3.4 An example of time-dependent learning rate
In this section we study a particular example of time-dependent learning rate, given by $(\eta_{r},\eta_{\bf u})=(r^{2},1)$. Note that instead of a constant in time, $\eta_{r}=r^{2}(t)$ is a function depends on time. In this case the dynamics is greatly simplified and is similar to gradient flow without normalization (1.11). Such simplification allows us to analyze the dynamics based on established methods, such as the argument with Bregman divergence in [8], and completely bypass the need of invariants and uniqueness results proved in Section 3.2.
However, gradient flow under this particular choice of learning rate ($(\eta_{r},\eta_{\bf u})=(r^{2},1)$) no longer exhibits the magnification effect as in the constant rate case ($(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$), and hence does not yield better bounds than the ones in previous works (1.8). It is nevertheless remarkable that the dynamics with certain choices of learning rate can be so different from the one with time-independent learning rate.
Lemma 3.14 (Dynamics reduction).
Suppose $r$ and ${\bf u}$ follow the gradient flow in (1.13) and (1.14) with $\|{\bf u}(0)\|=1$. Let ${\bf x}:=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Then
$\partial_{t}{\bf x}=-\left(\eta_{r}\frac{{\bf x}{\bf x}^{\top}}{\|{\bf x}\|_{2}^{2}}+\eta_{\bf u}(\|{\bf x}\|_{2}^{2}I-{\bf x}{\bf x}^{\top})\right)\nabla\mathcal{L}\left({\bf x}\right).$ (3.20)
Proof.
See Appendix A. ∎
Observe that if $(\eta_{r},\eta_{\bf u})=(r^{2},1)$, then (3.20) is greatly simplified as stated next.
Lemma 3.15 (Dynamic reduction, time-dependent learning rate).
$\partial_{t}{\bf x}=-\|{\bf x}\|_{2}^{2}\cdot\nabla\mathcal{L}\left({\bf x}\right).$ (3.21)
Proof.
Apply Lemma 3.14 with $(\eta_{r},\eta_{\bf u})=(r^{2},1)$ in (3.20). ∎
Note that without the additional factor $\|{\bf x}\|_{2}^{2}$, we are back to the setting that does not include weight normalization at all (1.3). Although the extra factor changes in time, since it is only a scalar as opposed to a vector or a matrix, it is possible to apply the same proof strategy as for gradient flow without normalization as in [8].
Definition 3.16 (Bregman Divergence).
$D_{F}(p,q)=F(p)-F(q)-\langle\nabla F(q),p-q\rangle.$ (3.22)
Lemma 3.17 ([6]).
Lemma 3.18 (Non-increasing Bregman Divergence).
$\partial_{t}D_{F}({\bf z},\tilde{{\bf x}})=-2L\|{\bf x}\|_{2}^{2}\cdot\mathcal{L}({\bf x})$ (3.23)
$F(\tilde{{\bf x}})=\begin{cases}\frac{1}{2}\langle\tilde{{\bf x}}\odot\log(\tilde{{\bf x}})-\tilde{{\bf x}},\bm{1}\rangle&\text{if }L=2\\ \frac{L}{2(2-L)}\langle\tilde{{\bf x}}^{\odot\frac{2}{L}},\bm{1}\rangle&\text{if }L>2.\end{cases}$ (3.24)
Proof.
This directly comes from Lemma 3.15 and Definition 3.16. ∎
Lemma 3.19 (Convergence, time-dependent learning rate).
Let $L\in\mathbb{N}$, $L\geq 2$, $(\eta_{r},\eta_{\bf u})=(r^{2},1)$, $\|{\bf u}(0)\|=1$ and $r(0)>0$. Denote ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Suppose that ${\bf b}$ is not identically zero and the solution set $S_{+}=\{{\bf z}\geq 0:{\bf A}{\bf z}={\bf b}\}$ is non-empty. Then $\lim_{t\to\infty}\mathcal{L}({\bf x}(t))=0$.
Proof.
See Appendix A. ∎
Proof of Theorem 2.4.
4 Experiments
In this section we test our method across different number of layers $L$, learning rate ratio ${\tilde{\eta}}$, and initialization scale
$\alpha:=\|(r_{0}{\bf u}_{0})^{\odot L}\|_{1}=\|\tilde{{\bf x}}_{0}\|_{1}.$ (4.1)
We will focus two things: the comparison of the reconstruction error between GD and WN-GD (GD with weight normalization), and how the learning rate ratio ${\tilde{\eta}}$ affects the reconstruction error.
We set the ambient dimension to be $N=1000$ and $M=150$. The matrix ${\bf A}$ is generated as
${\bf A}=\frac{1}{\sqrt{M}}{\bf G},\quad{\bf b}={\bf A}{\bf x}^{*}$
where ${\bf x}^{*}$ is the ground truth and ${\bf G}\in\mathbb{R}^{M\times N}$ has independent and standard normal distributed entries. Fix $s=10$. We examine the case where ${\bf x}^{*}\geq 0$ is $s$-sparse and has $\ell_{1}$-norm equals to $s$. All experiments are conducted with constant small step size. Each data point is an average over ten instances of random data and random initialization ${\bf x}_{0}$.
The reconstruction is defined as
$\varepsilon_{1}:=\|\tilde{{\bf x}}_{\infty}-{\bf x}^{*}\|_{1}.$
By Theorem 2.2, the difference $\|\tilde{{\bf x}}_{\infty}\|_{1}-\min_{{\bf z}\in S_{+}}\|{\bf z}\|_{1}$ should be small for small initialization. Due to the restricted isometry property of ${\bf A}$, the reconstruction error should also decrease as the initialization decreases.
In the first experiment we compare GD and WN-GD with fixed $L=2$ and ${\tilde{\eta}}=0.1$. In the second experiment we record the reconstruction error for different ${\tilde{\eta}}$ with fixed number of layer $L=2$ and initialization $\alpha=1$. In the third experiment we examine the performance of WN-GD for $L=2$ and $L=3$ with fixed ${\tilde{\eta}}=0.1$. In the forth experiment we no longer require ${\bf x}^{*}$ to be non-negative, and perform GD and WN-GD according to the loss function (2.8).
4.1 Compare GD and WN-GD
(a) WN-GD achieves smaller reconstruction error.
(b) Training loss of WN-GD for different initialization.
Figure 2: WN-GD yields significantly much smaller error than GD. The training loss converges to values close to zero.
The goal of the first experiment (Figure 2) is to compare GD and WN-GD among different initialization. We fix $L=2$ and ${\tilde{\eta}}=0.1$. In Figure 2(a), we compare the error $\varepsilon_{1}$ produced by the two algorithms. According to Theorem 2.2, for any fixed initialization satisfying the conditions in Theorem 2.2, WN-GD should yield smaller $\varepsilon_{1}$. In fact, the difference is quite significant.
In Figure 2(b) we observe that during training all loss decreases monotonically to vicinity of zero. According to experiments, convergence usually holds as long as the step size is sufficiently small and the initialization is smaller than the norm of the limit.
4.2 Compare learning rate ratio
(a) Smaller ${\tilde{\eta}}$ leaders to smaller error (for WN-GD).
(b) Training loss of WN-GD for different learning rate ratio.
Figure 3: As the learning rate ratio ${\tilde{\eta}}$ decreases, the error decreases. Note that when ${\tilde{\eta}}=10$, the assumption of Theorem 2.2 is violated, and we see that WN-GD is not better than GD.
In the second experiment (Figure 3) we aim to understand how ${\tilde{\eta}}$ affects the reconstruction error. Here we fix $L=2$ and $\alpha=1$. In Figure 3(a) we again analyze the error $\varepsilon_{1}$ among different algorithms. We see that in general as ${\tilde{\eta}}$ decreases, the error $\varepsilon_{1}$ decreases. In particular, such error is significantly smaller than the error of GD, which is represented by the horizontal line. Note that when ${\tilde{\eta}}$ is too large such that the condition of Theorem 2.2 is violated, WN-GD is no longer guaranteed to outperform GD.
4.3 Effects of Layer
In this section we compare results for $L=2$ and $L=3$. To ensure a fair comparison, we generate initialization in the following way. Choose an initialization scale $\alpha$ and the vector ${\bf u}_{0}={\bf 1}/\sqrt{N}$ so that $\|{\bf u}_{0}\|_{2}=1$. For each $L$, to make $\|(r_{0}{\bf u}_{0})^{\odot L}\|_{1}=\alpha$ we set
$r_{0}=\frac{\alpha^{\frac{1}{L}}}{\|{\bf u}_{0}\|_{L}}=\alpha^{\frac{1}{L}}N^{\frac{1}{2}-\frac{1}{L}}.$
The results are shown in Figure 4. In both cases ($L=2$ and $L=3$), we observe significant improvement of reconstruction error with weight normalization. Note that the significant improvement for $L=3$ requires smaller initialization than the case of $L=2$. Here the improvement ratio (reconstruction error for GD divided by reconstruction error for WN-GD) at $\alpha=0.1$ is $173810$ for $L=2$, and $92164$ for $L=3$.
Figure 4: Comparison between $L=2$ and $L=3$. WN-GD is better in both cases, but $L=3$ requires smaller initialization.
4.4 Sparse ground truth with positive and negative entries
(a) WN-GD still achieves smaller reconstruction error, but it is less significant than the positive case.
(b) Training loss of WN-GD are also decreasing, and the convergence rate seems to related to the initialization.
The setting of the forth experiment (Figure 5) is the same as the first experiment, except that the ground truth is not constrained to have positive entries, and the loss function we use here is the same as in (2.8)
5 Summary and Discussion
There are still many remaining questions such as
  1. 1.
  2. 2.
  3. 3.
    Can we extend the results here from gradient flow to gradient descent?
  4. 4.
Acknowledgement
R. Ward is grateful for support from AFOSR MURI FA9550-19-1-0005, NSF DMS 1952735, NSF HDR 1934932, HDR TRIPODS Phase II grant 2217058, and NSF CCF 2019844 H.C. and H.R. acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG) through the Collaborative Research Center Sparsity and Singular Structures (SFB 1481).
Data Availability Statement
The data underlying this article will be shared on reasonable request to the corresponding author.
Appendix A Appendix
Proof of Lemma 3.1.
First note that for
$r^{({a})}=r\quad\text{and}\quad{\bf u}^{({a})}={a}{\bf u}$ (A.1)
we obtain
Consequentially,
$\partial_{t}r^{({a})}$ $=-\eta_{r}\nabla_{r^{({a})}}\widetilde{\mathcal{L}}(r^{({a})},{\bf u}^{({a})})=\partial_{t}r,$ (A.2)
$\partial_{t}{\bf u}^{({a})}$ $=-{a}^{2}\eta_{\bf u}\nabla_{{\bf u}^{({a})}}\widetilde{\mathcal{L}}(r^{({a})},{\bf u}^{({a})})={a}\partial_{t}{\bf u}.$ (A.3)
Since (A.1) holds for $t=0$ and is preserved due to (A.2) and (A.3), (A.1) holds for all $t\geq 0$. ∎
Proof of Lemma 3.2.
A direct computation yields
because ${\bf u}^{\top}(I-\mathcal{P}_{{\bf u}})=0$. This implies the claim. ∎
Proof of Lemma 3.3.
Note that $\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})$ is local Lipschitz continuous in $r$, and $\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u})$ is local Lipschitz continuous in ${\bf u}$. Hence by the Picard–Lindelöf theorem the trajectory is unique. In particular, $\partial_{t}r$ and $u_{n}$ cannot reach zero at finite time, since this would otherwise contradict to the uniqueness of the trajectory (if we apply Picard–Lindelöf theorem backward in time). We now show that $r=0$ implies that $\partial_{t}r=0$ and $u_{n}=0$ implies that $\partial_{t}u_{n}=0$. By (3.1), (3.2), and (3.3), we have
$\partial_{t}r$ $=-\eta_{r}\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})=-\frac{\eta_{r}}{\|{\bf u}\|_{2}}{\bf u}^{\top}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)$ (A.4)
$=-\eta_{\bf u}\frac{r}{\|{\bf u}\|_{2}}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)-\eta_{\bf u}\frac{r\partial_{t}r}{\eta_{r}\|{\bf u}\|_{2}^{2}}{\bf u}.$ (A.5)
Proof of Lemma 3.5.
By chain rule we have
This completes the proof. ∎
Proof of Lemma 3.6.
By Lemma 3.2, $\|{\bf u}\|_{2}=\|{\bf u}_{0}\|_{2}=1$. By Lemma 3.5, we have
Note that by (3.2) and (3.3),
On the other hand,
Putting together the estimates, we obtain
By Gronwall’s inequality we get the linear convergence rate
This completes the proof. ∎
Proof of Lemma 3.10.
@@ Appendix A Appendix
Since $\gamma$ defined in (3.16) is both bounded above and below, it suffices to prove the case for ${\bf x}$. We will prove the statement by contradiction. Suppose $x_{j}$ is not bounded away from zero. Then, there exists a sequence $\{t_{k}\}_{k\in\mathbb{N}}$ such that $\lim_{k\to\infty}x_{j}(t_{k})=0$. By assumption there exists ${\bf v}>0$ such that ${\bf A}{\bf v}=0$. Then for any ${\bf z}$ such that $(I-\mathcal{P}_{{\bf A}}){\bf z}=0$, it holds
@@
Proof of Lemma 3.11.
@@
$\geq\bm{\Delta}\odot\log({\bf 1}+\bm{\Delta}\odot(\tilde{{\bf x}}^{(1)})^{\odot-1}).$ (A.6)
Since $\bm{\xi}\geq 0$ and $\langle\bm{\xi}(t),{\bf 1}\rangle$ converges to zero as $t\to\infty$, we can deduce that $\bm{\xi}$ converges to zero as $t\to\infty$. Together with (A.6) and the assumption that $\tilde{{\bf x}}^{(1)}$ is uniformly bounded above, we conclude that $\bm{\Delta}$ must also converges to zero. Since $\bm{\Delta}$ converges to zero, our conclusion follows.
@@
Proof of Lemma 3.12.
@@
$\|\mathcal{P}_{{\bf A}}(r{\bf u})^{\odot L}\|_{\infty}\leq B_{1}$ (A.7)
@@
$(I-\mathcal{P}_{{\bf A}})\cdot\begin{cases}\log(\widetilde{r}{\bf u})&\text{if}L=2\\ (\widetilde{r}{\bf u})^{\odot 2-L}&\text{if }L\neq 2\end{cases}$ (A.8)
@@
$\langle(r{\bf u})^{\odot 2},\log(\widetilde{r}{\bf u})\rangle\geq r^{2}\left(\log(\widetilde{r})-\frac{N}{2e}\right).$ (A.9)
@@
Note that $B_{1}\left\|\mathcal{P}_{{\bf A}}\log\left({\bf u}\right)\right\|_{1}$ is upper bounded by some constant $C$ because the entries of ${\bf u}$ are bounded both above and below. Combining this with (A.9), we have
@@
Proof of Lemma 3.14.
By (3.2) and (3.3),
@@
and hence (3.20) holds. ∎
Proof of Lemma 3.19.
@@
Let ${\bf z}\in S_{+}$. Since $r(0)>0$ and ${\bf u}(0)\geq 0$, we have ${\bf x}\geq 0$ for all $t\geq 0$ by Lemma 3.3. Hence $D_{F}({\bf z},\tilde{{\bf x}}(t))$ is well-defined. By Lemma 3.18, $D_{F}({\bf z},\tilde{{\bf x}}(t))$ is non-increasing in $t$ and hence bounded above by $D_{F}({\bf z},\tilde{{\bf x}}_{0})$. By Lemma 3.17, $D_{F}$ is non-negative and hence bounded below. Therefore by (3.23),
@@
Since $D_{F}({\bf z},\tilde{{\bf x}}_{0})<\infty$, by (A.10) $\|{\bf x}(t)\|_{2}$ cannot be bounded away from zero. This implies that there exists an increasing sequence $\{t_{k}\}_{k\in[\mathbb{N}]}$ such that $\lim_{k\to\infty}\|{\bf x}(t_{k})\|_{2}=0$. Together with the fact that ${\bf x}$ is non-negative, we get that for all $n\in[N]$, $\lim_{k\to\infty}x_{n}(t_{k})=0$, and consequently $\lim_{k\to\infty}\tilde{x}_{n}(t_{k})=0$.